# Video-XL: Extra-Long Vision Language Model for Hour-Scale Video Understanding

### v2

@@ Abstract
Although current Multi-modal Large Language Models (MLLMs) demonstrate promising results in video understanding, processing extremely long videos remains an ongoing challenge. Typically, MLLMs struggle with handling thousands of tokens that exceed the maximum context length of LLMs, and they experience reduced visual clarity due to token aggregation. Another challenge is the high computational cost stemming from the large number of video tokens. To tackle these issues, we propose Video-XL, an extra-long vision language model designed for efficient hour-scale video understanding. Specifically, we argue that LLMs can be adapted as effective visual condensers and introduce Visual Context Latent Summarization, which condenses visual contexts into highly compact forms. Extensive experiments demonstrate that our model achieves promising results on popular long video understanding benchmarks, despite being trained on limited image data. Moreover, Video-XL strikes a promising balance between efficiency and effectiveness, processing 1024 frames on a single 80GB GPU while achieving nearly 100% accuracy in the Needle-in-a-Haystack evaluation. We envision Video-XL becoming a valuable tool for long video applications such as video summarization, surveillance anomaly detection, and Ad placement identification.
Figure 1: Results on the Needle-in-a-haystack evaluation within a single 80G GPU. Compared to other video understanding models, Video-XL can achieve 100% accuracy within 1024 frames.
@@ 1 Introduction
In recent years, large language models (LLMs) [1, 2, 3, 4] have demonstrated exceptional capabilities in understanding, reasoning, and generating text across a wide range of open-ended tasks. Building on the broad applicability of LLMs, generative multimodal large language models (MLLMs) [5, 6, 7, 8, 9] have made significant progress in their ability to comprehend and generate content across multiple modalities, particularly images. Unlike static images, videos introduce a temporal dimension, consisting of sequences of frames that are crucial for understanding dynamic visual content in conjunction with textual input. Consequently, numerous studies [10, 11, 12, 13] have sought to adapt MLLMs to address the temporal complexities intrinsic to video sequences.
@@
Most existing models are optimized for short videos, typically under one minute, due to the inherent difficulty in establishing sufficient context for longer videos. To address this challenge, several approaches employ compression strategies, such as reducing the number of visual tokens [14, 15, 16] or introducing specialized memory mechanisms [17, 18] to extend the input visual context length. However, these methods inevitably lead to some degree of information loss during the compression process. Other studies attempt to extend the context length of the language backbone [19] or improve multi-node efficiency during the training process [20]. Nonetheless, they often overlook the computational costs incurred during the inference stage. For instance, LongVA and LongVila can process no more than 500 frames and 200 frames, respectively, on a single 80GB GPU.
@@
Expanding visual context length while maintaining computational efficiency remains a significant challenge. In response, we propose that the LLM itself can serve as an effective visual condenser with appropriate adaptations, leveraging its inherent context representation capabilities. By condensing the LLM’s raw activations (i.e., keys and values from the self-attention module) into highly compact forms, the LLM can perceive information from a broader context, even within a limited context window. While VoCo-LLaMA [21] attempted to distill visual tokens within the LLM, it struggles to handle long videos with thousands of tokens due to the LLM’s limited context window. To address this, we propose Visual Context Latent Summarization, which captures essential information from long visual contexts. Specifically, similar to  [22], we append visual summarization tokens (VST) to the original visual context, condense the raw activations, and process them in a streaming manner using a sliding window with an autoregressive approach. Moreover, to efficiently learn the condensed representation, we avoid introducing additional training strategies and focus solely on refining activation compression during the Supervised Fine-tuning (SFT) stage.
@@
Based on the aboved designs, we propose Video-XL, an extra-long vision language model for efficient hour-scale video understanding. Video-XL has the following highlights:
(i) Video-XL achieves promising performance on long video benchmarks—MLVU [23], Video-MME [24], LongVideoBench [25], and VNBench [26], despite being trained on limited image data.
@@
(ii) Video-XL strikes a promising balance between efficiency and effectiveness, processing 1024 frames on a single 80GB GPU while achieving nearly 100% accuracy in the Needle-in-a-Haystack evaluation.
(iii) Beyond general video understanding tasks, Video-XL shows potential for long video-specific tasks. For instance, it can provide summaries for hour-long movies, detect anomalous clips in surveillance videos, or identify inserted advertisements in long videos.
@@
Figure 2: Overview of Video-XL. Video-XL utilizes a unified visual encoding scheme for single images, multiple images, and videos. Visual tokens generated from the projector are first split into several chunks. The activations from the visual contexts in each chunk are then condensed into Visual Summarization Tokens (VSTs). Conditioned on the accumulated activations between chunks, Video-XL performs next-token prediction in an autoregressive manner.
@@ 2.1 Architecture and Pipeline
Large Language Model Backbone. We choose Qwen-2 [27] with 7B size as our LLM as it exhibits strong language understanding and reasoning capabilities.
Vision Encoder. We use CLIP-ViT-L [28] which is universally utilized to encode input images.
Cross-modality Projector. Following LLaVA [7], we leverage a two-layer MLP with a GELU activation function as a cross-modality projector to align the vision encoder and LLM.
@@ 2.2.2 Learning Method
$\boldsymbol{\Theta}$ denotes the parameters of the LLM, the vision encoder, the projector, the matrices for VSTs at each layer, $\boldsymbol{W}_{Q}^{v}$, $\boldsymbol{W}_{K}^{v}$, $\boldsymbol{W}_{V}^{v}$, and the token embedding of VST, $\boldsymbol{e}_{\langle\text{vs}\rangle}$ (we use one shared embedding for all VSTs). The standard auto-regression loss is minimized for training. Note that we exclude the beacon tokens from the training loss (setting their labels to -100) because they are only intended for compression. To flexibly support a diverse compression granularities, the compression ratio $\alpha_{i}$ for the $i$-th chunk is randomly sampled from $\{2,4,8\}$ during training. At inference, one can choose one compression ratio according to the specific efficiency requirement and stick to it for all chunks.
@@ 3 Training Strategy and Data
We employ a two-stage training strategy consisting of pre-training and fine-tuning. During the pre-training stage, we use the Laion-2M dataset [29] to optimize the projector, where visual embeddings from a pre-trained vision encoder are aligned with text embeddings from the LLM. In the fine-tuning stage, we apply visual instruction tuning to optimize both the projector and all LLM parameters. Additionally, we unfreeze the visual encoder to capture improved visual representations of temporal features. As discussed in Section 2.1, although Video-XL can process various multimodal datasets, we only use image data from Bunny 695k [29] in the training process for efficiency.
Model LLM-size MLVU VideoMME w/o sub. VideoMME w sub.
M-avg G-avg Short Medium Long Avg Short Medium Long Avg
@@
GPT-4V [1] - 49.2 5.35 70.5 55.8 53.5 59.9 73.2 59.7 56.9 63.3
GPT-4o [30] - 64.6 5.80 80.0 70.3 65.3 71.9 82.8 76.6 72.1 77.2
Gemini-1.5-Pro [31] - - - 81.7 74.3 67.4 75.0 84.5 81.0 77.4 81.3
Claude3-Opus [32] - 36.5 3.39 71.0 57.4 51.2 60.0 73.5 60.1 54.7 62.9
@@
VideoChat2 [33] 7B 47.9 3.99 48.3 37.0 33.2 39.5 52.8 39.4 39.2 43.8
LLaMA-VID [34] 7B 33.2 4.22 - - - - - - - -
VideoLLaVA [35] 7B 47.3 3.84 45.3 38.0 36.2 39.9 46.1 40.7 38.1 41.6
Shargpt4Video [36] 7B 46.4 3.77 48.3 36.3 35.0 39.9 53.6 39.3 37.9 43.6
InternVL-Chat-V1.5 [37] 20B 50.4 4.02 60.2 46.4 45.6 47.8 61.7 49.1 46.6 52.4
LongVA [19] 7B 56.3 4.33 61.1 50.4 46.2 52.6 61.6 53.6 47.6 54.3
LongVILA [20] 7B - - 61.8 49.7 39.7 50.5 - - - -
Long-LLaVA [15] 13B - - 61.9 51.4 45.4 52.9 66.2 54.7 50.3 57.1
Video-XL 7B 57.0 4.40 58.2 48.9 45.8 51.0 62.1 51.9 49.3 54.4
@@ 4.1 Benchmarks
We evaluated our methods on several mainstream video understanding benchmarks, including MLVU [23], Video-MME [24], VNBench [26] and LongVideoBench [25].
@@ 4.2.1 Long Video Benchmark Evaluation
MLVU is a long video understanding benchmark comprising 9 distinct tasks, categorized into Multiple-Choice and Generation tasks. As shown in Table 1, our model achieves leading results in both categories within the 7B model family. Furthermore, our approach surpasses some larger models like Intern-VL [37], showcasing a strong capability in long video understanding.
Video-MME is a comprehensive long video understanding benchmark that encompasses a highly diverse range of video types. It evaluates performance across short, medium, and long video lengths, as well as overall effectiveness. Trained only on image data, Video-XL achieves 51.0% and 54.4% accuracy without and with the assistance of subtitles, achieving competitive results with some state-of-the-art models (See Table 1).
VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. As shown in Table 2, Our model demonstrates leading results among open-source models on ordering and counting tass. Additionally, it surpasses GPT-4V in the final score.
LongVideoBench is a challenging benchmark designed to accurately retrieve and reason over detailed multimodal information from long inputs. With a context input of 256 frames, our model achieves a leading performance among all open-source models on the LongVideoBench-dev task, As demonstrated in Table 2,
Model LLM-size VNBench LongVideoBench
Retrieval Ordering Counting Average
@@
VideoChat2 [33] 7B 32.7 0.4 15.5 20.1 39.3
Video-LLaVA [35] 7B 23.8 1.1 12.4 12.4 39.1
LLaVA-NeXT-Video [38] 7B 44.2 0.4 15.5 20.1 43.5
ST-LLM [39] 7B 51.3 0.0 16.7 22.7 -
LongLLaVA [15] 13B 91.1 35.9 29.2 52.1 -
Video-XL 7B 72.0 46.3 33.5 50.6 48.8
@@ 4.2.2 Longer Context Evaluation
To demonstrate our model’s ability to process extremely long contexts, we conducted a needle-in-the-haystack test to evaluate its performance. As shown in Figure 1, both LLaVA-NexT-Video and LongLLaVA employ a naive position encoding extrapolation algorithm, yet they struggle to comprehend key information when provided with more input context. LongVA fine-tunes the LLM to handle extended input lengths, but it still faces high computational costs, limiting it to processing approximately 400 frames on a single 80G GPU. In contrast, our model achieved 100% accuracy with an 8×8$\times$8 × compression ratio and a 1024-frame input on the same hardware. This suggests that our model strikes an optimal balance between accuracy and computational efficiency.

### v3

@@ Abstract
Although current Multi-modal Large Language Models (MLLMs) demonstrate promising results in video understanding, processing extremely long videos remains an ongoing challenge. Typically, MLLMs struggle with handling thousands of visual tokens that exceed the maximum context length, and they suffer from the information decay due to token aggregation. Another challenge is the high computational cost stemming from the large number of video tokens. To tackle these issues, we propose Video-XL, an extra-long vision language model designed for efficient hour-scale video understanding. Specifically, we argue that LLMs can be adapted as effective visual condensers and propose Visual Context Latent Summarization which condenses visual contexts into highly compact forms. Extensive experiments demonstrate that our model achieves promising results on popular long video understanding benchmarks. For example, Video-XL outperforms the current state-of-the-art method on VNBench by nearly 10% in accuracy. Moreover, Video-XL presents an impressive balance between efficiency and effectiveness, processing 2048 frames on a single 80GB GPU while achieving nearly 95% accuracy in the Needle-in-a-Haystack evaluation.
Figure 1: (Left) The comparison of video MLLMs in performance and maximum frames. (Right) Results on the Needle-in-a-haystack evaluation within a single 80GB GPU. The x-axis represents the total number of frames in the video haystack. The y-axis shows the position where the needle image is located. Gray grids mean “Out of Memory”. Compared to other models, Video-XL can achieve nearly 95% accuracy with maximum 2048 frames.
@@ 1 Introduction
Inspired by the population of large language models (LLMs) [1, 2, 3], multimodal large language models (MLLMs) [4, 5, 6, 7] have made significant progress in their ability to comprehend and generate content across multiple modalities, particularly images. Beyond single image understanding, videos, consisting of sequence of frames, introduces the temporal dimension and brings significant challenges in methodology and computation.
@@
Existing MLLMs of video understanding divide videos to sequence of frames and adapt MLLMs to implement multi-image understanding. Due to the massive visual tokens of videos that exceed the context length of LLMs, popular methods [8, 9, 10, 11] can only process short videos, typically under one minute. To address this challenge, several methods employ compression strategies, such as reducing the number of visual tokens [12, 13, 14] and introducing specialized memory mechanisms [15, 16] to extend the input visual context length. However, these methods inevitably lead to certain degree of information loss. Other studies attempt to extend the context length of the language backbone [17] or improve multi-node efficiency during the training stage [18]. Nonetheless, they overlook the computational costs incurred during the inference stage.
@@
Expanding visual context length while maintaining computational efficiency remains a significant challenge in long video understanding. In this paper, we argue that long visual representations can be effectively compressed into more compact forms by leveraging the inherent context modeling capabilities of LLMs. To achieve this, we introduce Visual Context Latent Summarization. Specifically, we append Visual Summarization Tokens (VSTs) to the original visual context and condense the raw activations of visual latents (i.e., keys and values in the self-attention module) into highly compact representations. These condensed activations are processed in a streaming manner using a sliding window and an autoregressive approach. To efficiently learn this condensed representation, we refrain from introducing additional training strategies and instead focus on refining activation compression during the Supervised Fine-tuning (SFT) stage.
To enhance the model’s generalization ability for long video understanding, we develop a unified visual encoding scheme, which represents single images, multiple images and videos into a unified space. Additionally, we introduce a novel long video dataset called Visual Clue Ordering (VICO). Unlike most existing video datasets, which are typically short (less than 10 seconds) and feature static content with minimal changes between frames, VICO is specifically designed to handle longer videos and dynamic visual contexts. This dataset has three key features: (1) videos from diverse, open-domain categories, including movies, egocentric videos, games, and sports; (2) video lengths ranging from 2 to 10 minutes, providing a sufficient number of frames and substantial plot variations; (3) VICO encourages multi-detail retrieval and model understanding, offering a more comprehensive challenge than traditional captioning or question-answering tasks.
@@
Based on the above designs, we propose Video-XL, an extra-long vision language model for efficient hour-scale video understanding. Video-XL has the following highlights:
(i) Video-XL achieves promising performance on long video benchmarks. Specifically, Video-XL outperforms the current state-of-the-art method on VNBench [19] by nearly 10% in accuracy, as well as achieving leading performance in MLVU [20], Video-MME [21], LongVideoBench [22].
@@
(ii) Video-XL strikes a promising balance between efficiency and effectiveness, as shwon in Figure 1, processing 2048 frames on a single A100-80GB GPU while achieving nearly 100% accuracy in the Needle-in-a-Haystack evaluation.
(iii) Beyond general video understanding tasks, Video-XL shows potential for long video-specific tasks such as movie summarization, surveillance anomaly detection, and ad placement identification.
@@
Figure 2: Overview of Video-XL. Video-XL utilizes a unified visual encoding scheme for single images, multiple images, and videos. Visual tokens generated from the projector are first split into several chunks. The activations from the visual contexts in each chunk are then condensed into Visual Summarization Tokens (VSTs). Conditioned on the accumulated activations between chunks, Video-XL performs next-token prediction in an autoregressive manner. For simplicity, we only present the cases with compression ratios of 2, 4, and 8.
@@ 2.1 Architecture and Pipeline
We selected Qwen-2 [23] with 7B parameters as our LLM due to its strong language understanding and reasoning capabilities. For the Vision Encoder, we opted for CLIP-ViT-L [24] instead of a dedicated video encoder. This choice was motivated by CLIP’s strong generalization capabilities in various visual understanding tasks, as well as its ability to accept an arbitrary number of frames, whereas some video encoders are limited by the number of input frames. Following LLaVA [6], we leverage a two-layer MLP with a GELU activation function as a cross-modality projector to align the Vision Encoder and LLM. Video-XL has a flexible combination of the three modules which will be introduced in the future.
@@ 2.2 Visual Context Latent Summarization
$\boldsymbol{\Theta}$ denotes the parameters of the LLM, the vision encoder, the projector, the matrices for VSTs at each layer, $\boldsymbol{W}_{Q}^{v}$, $\boldsymbol{W}_{K}^{v}$, $\boldsymbol{W}_{V}^{v}$, and the token embedding of VST, $\boldsymbol{e}_{\langle\text{vs}\rangle}$ (we use one shared embedding for all VSTs). The standard auto-regression loss is minimized for training. Note that we exclude the VSTs from the training loss (setting their labels to -100) because they are only intended for compression. To flexibly support a diverse compression granularities, the compression ratio $\alpha_{i}$ for the $i$-th chunk is randomly sampled from $\{2,4,8,12,16\}$ during training. At inference, one can choose one compression ratio according to the specific efficiency requirement and stick to it for all chunks.
@@ 3 Training Strategy and Data
We employ a two-stage training strategy consisting of pre-training and fine-tuning. During the pre-training stage, we use the Laion-2M dataset [26] to optimize the projector, where visual embeddings from a pre-trained vision encoder are aligned with text embeddings from the LLM. In the fine-tuning stage, we apply visual instruction tuning to optimize both the projector and all LLM parameters. Additionally, we unfreeze the visual encoder to capture improved visual representations of temporal features. As discussed in Section 2.1, we fully leverage the capabilities of the MLLM across various multimodal datasets. For single-image data, we use the Bunny 695k dataset [26] and 57k images from Sharegpt-4o [27]. For multi-image data, we utilize 5k data points extracted from MMDU [28]. For video data, we collect videos of varying durations, including 32k samples from NExT-QA [29] 2k video samples from Sharegpt-4o [27], 10k samples from CinePile [30], and 11k in-house samples with GPT-4V annotations for video captions. Additionally, we curate 10k long video samples called VICO, which will be discussed in the next section.
3.1 Video Clue Order Dataset
To enhance long video understanding and unlock the potential of visual compression, we developed an automated long-video data production pipeline alongside a high-quality dataset called VIsual Clue Order (VICO). The pipeline, begins by sourcing long videos from CinePile or video platforms such as YouTube, covering open-domain content including movies, documentaries, games, sports, and more. Each long video is segmented into 14-second clips. For each clip, we employ the VILA-1.5 model [31] to generate detailed descriptions. Based on these captions, we utilize ChatGPT to arrange the clues in temporal order. VICO challenges models to retrieve key frames and detect temporal changes, thus improving their long video comprehension capabilities.
Model Size MLVU dev MLVU test VideoMME w/o sub. VideoMME w sub.
M-avg G-avg M-avg G-avg Short Medium Long Avg Short Medium Long Avg
@@ 3.1 Video Clue Order Dataset
GPT-4V [1] - 49.2 5.35 43.3 4.67 70.5 55.8 53.5 59.9 73.2 59.7 56.9 63.3
GPT-4o [32] - 64.6 5.80 54.9 5.87 80.0 70.3 65.3 71.9 82.8 76.6 72.1 77.2
Gemini-1.5-Pro [33] - - - - - 81.7 74.3 67.4 75.0 84.5 81.0 77.4 81.3
Claude3-Opus [34] - 36.5 3.39 25.0 3.25 71.0 57.4 51.2 60.0 73.5 60.1 54.7 62.9
@@
VideoChat2 [35] 7B 47.9 3.99 35.1 3.99 48.3 37.0 33.2 39.5 52.8 39.4 39.2 43.8
LLaMA-VID [36] 7B 33.2 4.22 17.2 3.43 - - - - - - - -
VideoLLaVA [37] 7B 47.3 3.84 30.7 3.68 45.3 38.0 36.2 39.9 46.1 40.7 38.1 41.6
Shargpt4Video [38] 7B 46.4 3.77 33.8 3.63 48.3 36.3 35.0 39.9 53.6 39.3 37.9 43.6
InternVL-Chat-V1.5 [39] 20B 50.4 4.02 37.3 3.96 60.2 46.4 45.6 47.8 61.7 49.1 46.6 52.4
LongVA [17] 7B 56.3 4.33 41.1 3.91 61.1 50.4 46.2 52.6 61.6 53.6 47.6 54.3
LongVILA [18] 7B - - - - 61.8 49.7 39.7 50.5 - - - -
Video-CCAM [40] 14B 63.1 4.01 42.9 3.57 62.2 50.6 46.7 53.2 66.0 56.3 49.9 57.4
Long-LLaVA [13] 7B - - - - 61.9 51.4 45.4 52.9 66.2 54.7 50.3 57.1
Video-XL 7B 64.9 4.50 45.5 4.21 64.0 53.2 49.2 55.5 67.4 60.7 54.9 61.0
@@ 4.1 Benchmarks
We evaluated our methods on several mainstream video understanding benchmarks. To assess long video capabilities, we used MLVU [20], VNBench [19], Video-MME [21] and LongVideoBench [22]. For short video evaluation, we tested on MVbench [35] and NExT-QA [29].
@@ 4.2.1 Long Video Benchmark Evaluation
MLVU is a long video understanding benchmark comprising a dev set and test set, categorized into Multiple-Choice and Generation tasks. Compared to the dev set, the test set is more challenging which includes harder tasks and longer videos. As shown in Table 1, our model achieves leading results in both tasks. It is also worth noting that Video-XL even surpasses GPT-4o in the dev task, showcasing a strong capability in long video understanding.
Video-MME is a comprehensive long video understanding benchmark that encompasses a highly diverse range of video types. It evaluates performance across short, medium, and long video lengths, as well as overall effectiveness. Video-XL achieves 55.5% and 61.0% accuracy without and with the assistance of subtitles, achieving competitive results with some state-of-the-art models (See Table 1).
VNBench is a synthetic benchmark designed to evaluate models’ long-context abilities, covering tasks such as retrieval, ordering, and counting. As shown in Table 2, Video-XL demonstrates leading results among open-source models (outperforms nearly 10 % in accuracy than the state-of-the-art model.) Additionally, it surpasses GPT-4V and achieves competitive results with GPT-4o in the average score.
LongVideoBench is a challenging benchmark designed to accurately retrieve and reason over detailed multimodal information from long inputs. With a context input of 256 frames, our model achieves a leading performance among all open-source models within 7B sizes on the dev and test task, as demonstrated in Table 2.
Model Size VNBench LongVideoBench
Retrieval Ordering Counting Average dev test
@@
VideoChat2 [35] 7B 32.7 0.4 15.5 20.1 39.3 41.2
Video-LLaVA [37] 7B 23.8 1.1 12.4 12.4 39.1 37.6
LLaVA-NeXT-Video [41] 7B 44.2 0.4 15.5 20.1 43.5 43.5
ST-LLM [42] 7B 51.3 0.0 16.7 22.7 - -
LongLLaVA [13] 7B 91.1 35.9 29.2 52.1 - -
Video-XL 7B 80.0 80.6 24.0 61.6 49.5 51.3
@@ 4.2.2 Longer Context Evaluation
To demonstrate our model’s ability to process extremely long contexts, we conducted a needle-in-the-haystack test to evaluate its performance. As shown in Figure 1, both LLaVA-NexT-Video and LongLLaVA employ a naive position encoding extrapolation algorithm, yet they struggle to comprehend key information when provided with more input context. LongVA fine-tunes the LLM to handle extended input lengths, but it still faces high computational costs, limiting it to processing approximately 400 frames on a single 80G GPU. In contrast, our model achieved nearly 95% accuracy with an 16×16$\times$16 × compression ratio and a 2048-frame input on the same hardware. This suggests that our model strikes an optimal balance between accuracy and computational efficiency.
Model Size NeXT-QA MVBench
Proprietary Models
GPT-4V [1] - - 43.5
GPT-4o [32] - 76.0 -
Open-source MLLMs
SeViLA [43] 4B 73.8 -
mPLUG-Owl [44] 7B 33.8 29.7
Video-LLaVA [37] 7B 40.2 -
VideoChat2 [35] 7B 78.6 51.9
LLaVA-NeXT-Video [41] 7B 41.6 61.3
TimeChat [45] 7B - 38.5
ST-LLM [42] 7B - 54.9
PLLaVA [46] 7B 45.6 58.1
LongLLaVA [13] 7B - 54.6
VideoLLaMB [10] 7B 71.1 52.5
Video-XL* 7B 77.2 55.3
Table 3: Results on Next-QA and MVBench.
4.2.3 Other Video Benchmarks Evaluation
MVBench is a comprehensive benchmark with average about 16s videos length, including 20 distinct video tasks, each with 200 questions designed to probe the model’s understanding of video content. Although our model is tailed for long video understanding, it also demonstrates promising results when dealing with short videos.
NExT-QA aims to comprehensively benchmark tempoal and casual reasoning of models. The average video lengths is 45 seconds, thus, we utilize 32 frames for testing. It can be seen that out models demonstrating astonishing results, achieving state-of-the-arts results.